OpenStudy (anonymous):

The slope of the graph of the linear function parallel to 6x-2y-4=0

OpenStudy (anonymous):

is it clear that what you are really being asked is to find the slope of \(6x-2y-4=0\) ?

OpenStudy (anonymous):

I'm confused to this question

OpenStudy (anonymous):

yes, i can see why, because it is worded both poorly and incorrectly

OpenStudy (anonymous):

let me rephrase it find the slope of a line parallel to the line given by the equation \(6x-2y-4=0\) is that better?

OpenStudy (anonymous):

"no" is a fine answer, i am just asking if it is clear now what the question wants

OpenStudy (anonymous):

yes that's better. Sorry late reply my computer is kinda laggy

OpenStudy (anonymous):

ok then what you want is to find the slope of \[6x-2y-4=0\] that will be your answer do you know how to find it?

OpenStudy (anonymous):

is it possible to get the parallel function of 6x-2y-4=0?

OpenStudy (anonymous):

no the slope is just the slope there are an infinite number of lines parallel to \(6x-2y-4=0\)

OpenStudy (anonymous):

what did you get for the slope?

OpenStudy (anonymous):

sorry for the bother i know how to get the slope . I was just confused of this question

OpenStudy (anonymous):

ok hope you got 3